OpenStudy (anonymous):

Solve the compound inequality 9 + 3n 6 or 5n < 25.

OpenStudy (anonymous):

9 + 3n greater or equal to 6 or 5n < 25.

OpenStudy (anonymous):

can you simplify them?

OpenStudy (anonymous):

idk how to

OpenStudy (anonymous):

pretend they are equalities, can you simplify 9 + 3n = 6 ?

OpenStudy (anonymous):

12n=6??

OpenStudy (anonymous):

add -9 to both sides

OpenStudy (anonymous):

oh you should of said like an equation.

OpenStudy (anonymous):

15

OpenStudy (anonymous):

3n=15

OpenStudy (anonymous):

watch the sign (-) 9 + 3n = 6 so add (-9) to both sides

OpenStudy (anonymous):

3n=6 - 9

OpenStudy (anonymous):

yes, thats it... now what is 6 - 9 ?

OpenStudy (anonymous):

-3

OpenStudy (anonymous):

yes, so 3n = -3 and when you divide both sides by 3 what do you get?

OpenStudy (anonymous):

n=-1

OpenStudy (anonymous):

great, now that first equation makes sense... n >= 1 so let's work on the second one 5n < 25 any ideas?

OpenStudy (anonymous):

pretend the < is an = for this part

OpenStudy (anonymous):

n=5

OpenStudy (anonymous):

exactly!!! and that is how you handle these... but be careful if you divide by a negative

OpenStudy (anonymous):

when you divide both sides by a negative number, you switch the sign

OpenStudy (anonymous):

but we didnt have to do that here.... so we have n >= 1 and n < 5

OpenStudy (anonymous):

ok thank you so much give me a medal plzz

OpenStudy (anonymous):

and that's all there is to it (unless you have to graph it or something